A solenoid has radius 2.20 cm, length 0.90 m and 1000 total number of turns. a) Determine the magnetic field energy density inside the solenoid. b) Calculate the total energy stored in the magnetic field inside the solenoid. Neglect the end effects. 


1
Expert's answer
2022-01-18T09:34:34-0500

a)

"\omega=\frac{\mu_0n^2I^2}2=0.23~\frac J{m^3},"

b)

"W=\omega V=\omega \pi r^2l=0.1~mJ."